# FAST drift scan survey for Hi intensity mapping: II. stacking-based primary beam construction for FAST L-band 19 feeds at $1.4$ GHz

### v1

@@ 3.3 Source selection criteria
The matrix $\bm{P}$ represents the projection between the maps and the time-ordered domains. Thus, it has the shape of $t\times p$, where $p$ is the number of pixels of the sky model and $t$ is the number of time stamps. Utilizing the complete time-ordered observational data yields a large $\bm{P}$ matrix, rendering the subsequent matrix operation both time- and memory-intensive. On the other hand, not all the observational data are useful for beam measurements. For instance, the RFI-contaminated data are removed at the first stage of data selection. Although we chose a frequency band that is relatively RFI-free, there still are time stamps that are badly contaminated by transient objects, e.g. the navigation satellites. In addition, the sources with other sources around, are not the ideal calibrator for characterizing the beam pattern. Thus, we only chose the time-ordered data with bright isolated sources for the following beam stacking analysis, which significantly reduces the amount of observation data.
@@
where $S_{i}$ is the flux density of the source around and $S_{j}$ is the flux density of the central source, $(\delta\alpha\cos(\beta),\delta\beta)$ are the separation angle between the two sources in the R.A. and decl. directions. The PSF is assumed to be a Gaussian kernel with $\sigma_{\rm PSF}=\theta_{\rm FWHM}/2\sqrt{2\ln 2}$, where $\theta_{\rm FWHM}=2.9^{\prime}$ is the Full Width at Half Maximum of FAST prior beam. We set a threshold of $\epsilon=0.05$, i.e. the sources are removed if there is more than one nearby source’s point-spread flux density over $5\%$ of target source flux density.
@@ 4 Results and Discussion
The positions of the 19 feeds within the FAST L-band 19-feed array (19FA) frame are shown in Figure 2. Here, the $H$-axis and $V$-axis correspond to the horizontal and vertical axes. We group the 19 feeds into four subsets according to their separation distance to the feed array center, i.e. as shown in Figure 2, the center beam (red), inner-circle beams (blue), middle-circle beams (green), and the outer-circle beams (black). During the drift scan observation, the feed array rotated by $23.4^{\circ}$. The cyan dashed line with an arrow indicates the drift scanning direction, which is aligned with the R.A. direction in the observational frame.
@@ 4.1 The stacking-based beam pattern
Figure 3: The stacked beam patterns for four different beams (Beam01, Beam02, Beam08, and Beam19) obtained from the data between 1375-1425MHz. The four columns represent the data from three different sources: FATHOMER (first column), CRAFTS (second column), and a combination of both (third column). Each row corresponds to a unique beam, with each beam having a distinct distance from the central beam center.
@@
Figure 4: The stacked beam patterns for four different beams (Beam01, Beam02, Beam08, and Beam19) obtained from the data between 1375-1425MHz. The four columns represent the polarization results for XX (vertical), YY (horizontal), and Stokes I (total intensity), respectively. Each row corresponds to a unique beam, with each beam having a distinct distance from the central beam center.
@@
To investigate the consistency of the stacking-based beam pattern reconstruction method, we select observations with similar decl. coverage from FATHOMER and CRAFTS. We then perform the stacking analysis separately on each dataset as well as on the combined dataset. We combine the two linear polarizations, i.e. XX and YY, to increase the signal-to-noise ratio. The reconstructed beam patterns are shown in Figure 3, where the results of FATHOMER, CRAFTS, and their combined dataset are shown in the left, middle, and right columns, respectively. From the top to the bottom rows, the results are Beam01, Beam02, Beam19, and Beam08, which are examples of beams located in the center, inner circle, middle circle, and the outer circle of the FAST L-band feed array, respectively. All the beam patterns are shown in the observation frame, where the horizontal and vertical axes are aligned with the R.A. and decl. direction.
@@
The shape of the main beam can be constructed by stacking either with the FATHOMER or CRAFTS dataset. The side lobes of Beam02, Beam19, and Beam08 are also visible with each dataset. However, significant data deficiency, which is due to either lack of observation or RFI flagging, is also visible, especially within the size of the main beam. Such data deficiencies are substantially eliminated by combining the two datasets, as shown in the right columns, where the blank pixels within the main beam are populated with supplementary datasets. The remaining blank pixels outside the main beam are due to the observation noise.
@@
The shapes of the main beam and side lobes show distinct variations across different beam positions. Beam01 displays a symmetric circular main beam profile, while Beam02, Beam08, and Beam19 exhibit more pronounced distortions, each showing an asymmetry with one side compressed and the other elongated. Notably, the side lobe shapes are the most striking feature among the four beams. Beam01’s side lobe is suppressed, whereas those of Beam02, Beam08, and Beam19 are visible. The side lobes of Beam08 and Beam19, in particular, are especially pronounced, with clearly discernible contours.
@@ 4.2 The beam pattern model
Figure 5: The residuals of the measured beam pattern and the fitting models. The top panels show the residuals for the Gaussian model while the bottom panels show the residuals for the ZP model. From the left to the right columns, there are the residuals for Beam01, Beam02, Beam19, and Beam08, which are examples of the subset of the center beam, inner-circle beams, middle-circle beams, and outer-circle beams.
@@
For the center beam, both models yield similar results, i.e., the residual distributions closely resembled the normal distribution. As long as the beam profile is circularly symmetric and the side lobes are suppressed, the Gaussian profile is still a good approximation of the beam pattern shape.
In contrast, the residual distributions of the off-center beams, e.g., the inner-circle beams, middle-circle beams, and outer-circle beams, reveal significant differences between the ZP model and the Gaussian model. Particularly, the Gaussian model has residual distributions that deviate from the normal distribution, significantly, which indicates the systematic differences between the Gaussian model and the beam pattern measurements. On the other hand, the ZP model residual behaves in the unique normal distribution for all the beam subsets. The results show that the ZP model successfully characterizes the primary features of the beam pattern for both the symmetric and asymmetric cases.
@@ 4.4 The noise-filtered beam pattern
The fitting coefficients of ZP modes with $|m|\geqslant 3$, as shown with the yellow markers in Figure 7, are mostly lower than $1\%$ of the maximum fitting coefficients. ZP modes with $|m|\geqslant 3$ generally characterize highly oscillated structures. Although such structures are tiny compared to the main beam profile, it is still important to be fully characterized for future Hi IM experiments. However, given current measurement uncertainties, such weak structures are mostly dominated by observational noise. The equipment’s thermal noise is one of the origins of measurement uncertainty. On the other hand, the flux density variation of the NVSS sources potentially impacts the measurement. Both the thermal noise and flux density variation are systematic and can be eliminated via a large amount of observation. In this work, we drop the ZP modes with fitting coefficients less than $1\%$ of the maximum fitting coefficients. Eventually, the number of ZP modes adopted varies across different beams. The central beam has $31$ ZP modes, which is the minimal number of ZP modes. The ZP mode numbers increase with the beam getting further from the feed array center, e.g., the inner-circle beams have $\sim 33$ ZP modes on average, and the middle- and outer-circle beams have $\sim 44$ on average, respectively. The noise-filtered beam patterns for all the 19 feeds are shown in Figure 9, where the relative positions of the panels correspond to the feed pointing direction in the sky. All the beam patterns are rotated to the feed coordinates. The yellow, green, and blue contours in each panel indicate the $50\%$, $10\%$, and $1\%$ of the maximum, respectively. As discussed in Section 4.1, we focus on the Stokes I component and all the beam patterns are averaged across the frequency range of $1375\,{\rm MHz}$$1425\,{\rm MHz}$.
@@
The beam patterns between the $10\%$ and the $1\%$ contours are identified as the side lobes. Significantly, the side lobe profiles vary across the 19 beams. After filtering out the noise components, the asymmetric side lobe profiles are clearer than those shown in Section 4.1. The side lobes are getting more pronounced with the larger separation distance between the beam and the feed array center. In addition, the direction of the side lobe distortion relates to the feed array’s relative position. Specifically, the side lobes become more pronounced on the side of the beam closer to the center of the feed array along the radial direction. Although the circular symmetry of the off-center beam is disrupted, the beam pattern still retains axial symmetry, with its axis of symmetry aligned with the radial direction of the feed array.
@@ 5 Conclusion
We develop a beam reconstruction method based on stacking analysis. This approach utilizes drift-scan observational data, adopting radio continuum point sources within the drift-scan field, such as NVSS sources, to reconstruct the antenna’s beam pattern. This method minimizes the impact of beam variations caused by changes in antenna configuration on actual observations, thus optimizing the accuracy of beam measurement.
We adopt the ZP modes as the analytic basis and decompose the two-dimensional beam pattern using these modes. ZP modes with fitting coefficients below $1\%$ of the maximum coefficient are discarded. Specifically, $31$ ZP modes are used to construct the central beam pattern; on average, approximately $33$ ZP modes are used for the inner-circle beams, and around $44$ ZP modes for the middle- and outer-circle beams.

### v2

@@ 3.3 Source selection criteria
The matrix $\bm{P}$ represents the projection between the maps and the time-ordered domains. Thus, it has the shape of $t\times p$, where $p$ is the number of pixels of the sky model and $t$ is the number of time stamps. Utilizing the complete time-ordered observational data yields a large $\bm{P}$ matrix, rendering the subsequent matrix operation both time- and memory-intensive. On the other hand, not all the observational data are useful for beam measurements. For instance, the RFI-contaminated data are removed at the first stage of data selection. Although we chose a frequency band that is relatively RFI-free, there still are time stamps that are badly contaminated by transient objects, e.g. the navigation satellites. In addition, calibrators with other sources nearby are not ideal for characterizing the beam pattern. Thus, we only chose the time-ordered data with bright isolated sources for the following beam stacking analysis, which significantly reduces the amount of observation data.
@@
where $S_{i}$ is the flux density of the neighboring source and $S_{j}$ is the flux density of the central source, $(\delta\alpha\cos(\beta),\delta\beta)$ are the separation angle between the two sources in the R.A. and declination directions. The PSF is assumed to be a Gaussian kernel with $\sigma_{\rm PSF}=\theta_{\rm FWHM}/2\sqrt{2\ln 2}$, where $\theta_{\rm FWHM}=2.9^{\prime}$ is the Full Width at Half Maximum of FAST prior beam. We set a threshold of $\epsilon=0.05$, i.e. the sources are removed if there is more than one nearby source’s point-spread flux density over $5\%$ of the target source flux density.
@@ 4 Results and Discussion
The positions of the 19 feeds within the FAST L-band 19-feed array (19FA) frame are shown in Figure 2. Here, the $H$-axis and $V$-axis correspond to the horizontal and vertical feed axes. We group the 19 feeds into four subsets according to their separation distance to the feed array center, i.e. as shown in Figure 2, the center beam (red), inner-circle beams (blue), middle-circle beams (green), and the outer-circle beams (black). During the drift scan observation, the rotation angle of the feed array was fixed at $23.4^{\circ}$. The cyan dashed line with an arrow indicates the drift scanning direction, which is aligned with the R.A. direction in the observational frame.
@@ 4.1 The stacking-based beam pattern
Figure 3: The stacked beam patterns for four different beams (Beam01, Beam02, Beam08, and Beam19) were obtained from the data between 1375-1425MHz. The three columns represent the data from three different sources: FATHOMER (first column), CRAFTS (second column), and a combination of both (third column). Each row corresponds to a unique beam, with each beam having a distinct distance from the central beam center.
@@
Figure 4: The stacked beam patterns for four different beams (Beam01, Beam02, Beam08, and Beam19) were obtained from the data between 1375-1425MHz. The three columns represent the polarization results for XX (vertical), YY (horizontal), and Stokes I (total intensity), respectively. Each row corresponds to a unique beam, with each beam having a distinct distance from the central beam center.
@@
To investigate the consistency of the stacking-based beam pattern reconstruction method, we select observations with similar declination coverage from FATHOMER and CRAFTS. We then perform the stacking analysis separately on each dataset as well as on the combined Stokes I dataset via Equation (5) to increase the signal-to-noise ratio. The reconstructed beam patterns are shown in Figure 3, where the results of FATHOMER, CRAFTS, and their combined dataset are shown in the left, middle, and right columns, respectively. From the top to the bottom rows, the results are Beam01, Beam02, Beam19, and Beam08, which are examples of beams located in the center, inner circle, middle circle, and the outer circle of the FAST L-band feed array, respectively. All the beam patterns are shown in the observation frame, where the horizontal and vertical axes are aligned with the R.A. and declination direction.
@@
The shape of the beam pattern can be constructed by stacking either with the FATHOMER or CRAFTS dataset. The circular patterns at the center are identified as the main lobe, while the asymmetric patterns offset from the center are referred to as side lobes. A detailed definition of the main lobe is provided in Section 4.6. The side lobes of Beam02, Beam19, and Beam08 are visible with each dataset. However, significant data deficiency, which is due to either lack of observation or RFI flagging, is also visible, even within the main lobe. Such data deficiencies are substantially eliminated by combining the two datasets, as shown in the right columns, where some blank pixels within the main lobe are filled by the complementary dataset. The remaining blank pixels, particularly those outside the main beam with negative power values, are likely due to observational noise, as TOD pre-whitening involves subtracting the temporal baseline.
@@
The shapes of the main beam and side lobes show coma distortions across different beam positions. Beam01 displays a symmetric circular main beam profile, while Beam02, Beam08, and Beam19 exhibit more pronounced distortions, each showing an asymmetry with one side compressed and the other elongated. Notably, the side lobe shapes are the most striking feature among the four beams. Beam01’s side lobe is suppressed, whereas those of Beam02, Beam08, and Beam19 are visible. The side lobes of Beam08 and Beam19, in particular, are especially pronounced, with clearly discernible contours.
@@ 4.2 The beam pattern model
Figure 5: The residuals of the measured beam pattern and the fitting models. These residuals are divided by the main lobe peak value of fitting models. The top panels show the residuals for the Gaussian model while the bottom panels show the residuals for the ZP model. From the left to the right columns, there are the residuals for Beam01, Beam02, Beam19, and Beam08, which are examples of the subset of the center beam, inner-circle beams, middle-circle beams, and outer-circle beams.
@@
For the center beam, both models yield similar results, i.e., the residual distributions closely resembled the normal distribution. As long as the beam profile is circularly symmetric and the coma lobes are suppressed, the Gaussian profile is still a good approximation of the beam pattern shape.
In contrast, the residual distributions of the off-center beams, e.g., the inner-circle beams, middle-circle beams, and outer-circle beams, reveal significant differences between the ZP model and the Gaussian model. Particularly, the Gaussian model has residual distributions that deviate from the normal distribution, significantly, which indicates the systematic differences between the Gaussian model and the beam pattern measurements. On the other hand, the ZP model residual behaves in the unique normal distribution for all the beam subsets. The results show that the ZP model successfully characterizes the primary features of the beam pattern for both the symmetric and asymmetric cases, as the ZP includes coma distortion terms.
@@ 4.4 The noise-filtered beam pattern
The fitting coefficients of ZP modes with $|m|\geqslant 3$, as shown with the yellow markers in Figure 7, are mostly lower than $1\%$ of the maximum fitting coefficients. ZP modes with $|m|\geqslant 3$ generally characterize highly oscillated structures. Although such structures are tiny compared to the main beam profile, it is still important to be fully characterized for future Hi IM experiments. However, given current measurement uncertainties, such weak structures are mostly dominated by observational noise. The equipment’s thermal noise is one of the origins of measurement uncertainty. On the other hand, the flux density variation of the NVSS sources potentially impacts the measurement. Both the thermal noise and flux density variation are systematic and can be eliminated via a large amount of observation. In this work, we drop the ZP modes with fitting coefficients less than $1\%$ of the maximum fitting coefficients. Eventually, the number of ZP modes adopted varies across different beams. The central beam has $31$ ZP modes, which is the minimal number of ZP modes. The ZP mode numbers increase with the beam getting further from the feed array center, e.g., the inner-circle beams have $\sim 33$ ZP modes on average, while the middle- and outer-circle beams have $\sim 44$ on average. The noise-filtered beam patterns for all the 19 feeds are shown in Figure 9, where the relative positions of the panels correspond to the feed pointing direction in the sky. All the beam patterns are rotated to the feed coordinates. The yellow, green, and blue contours in each panel indicate the $50\%$, $10\%$, and $1\%$ of the maximum, respectively. As discussed in Section 4.1, we focus on the Stokes I component and all the beam patterns are averaged across the frequency range of $1375\,{\rm MHz}$$1425\,{\rm MHz}$.
@@
The beam features between the $1\%$ and the $10\%$ contours are identified as the side lobes. Significantly, the side lobe profiles vary across the 19 beams. After filtering out the noise components, the asymmetric side lobe profiles are clearer than those shown in Section 4.1. The coma lobes are getting more pronounced with the larger separation distance between the beam and the feed array center. In addition, the direction of the side lobe distortion relates to the feed array’s relative position. Specifically, the coma lobes become more pronounced on the side of the beam closer to the center of the feed array along the radial direction. Although the circular symmetry of the off-center beam is disrupted, the beam pattern still retains axial symmetry, with its axis of symmetry aligned with the radial direction of the feed array.
@@ 5 Conclusion
We develop a beam reconstruction method based on stacking analysis. This approach utilizes drift-scan observational data, using radio continuum point sources within the drift-scan field, such as NVSS sources, to reconstruct the antenna’s beam pattern. This method minimizes the impact of beam variations caused by changes in antenna configuration on actual observations, thus optimizing the accuracy of beam measurement.
We adopt the ZP modes as the analytic basis and decompose the two-dimensional beam pattern using these modes. ZP modes with fitting coefficients below $1\%$ of the maximum coefficient are discarded to prevent overfitting. Specifically, $31$ ZP modes are used to construct the central beam pattern; on average, approximately $33$ ZP modes are used for the inner-circle beams, and around $44$ ZP modes for the middle- and outer-circle beams.